[Brahms-l] run status

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Sun Jan 16 2005 - 16:47:28 EST
A couple of problems was encountered this afternoon;

-- The TPC pedestals for tmp2 was bad for a rather large number of pad's particular in the last rows. This is likely due to the fact that no pedestals has been installed since Dec 14. The default icon as described in the 'manual' does not work/exists.
To do the job execute in a terminal window
>perl ~daq/scripts/tpcped.perl
and follow instructions.

-- The L1458A mainframe and Tiles HV went daed this morning;the circuit breaker in the IR was tripped and on of two powersupplies in the MF dead. Bob Scheetz restored the systems
and hopefully they will keep up until the Wednesday maintenance day.

-- The shhort term program should be
 -- MRS continue high field running
-- FS 4 deg
     1219A : 200 k tr2
      1219B: 400 k tr 2
     1723B  400k tr 2
     1723A  200k tr6

The down scale is considerable since the MCR has managed to get initial ZDC rates above 10K.

--Some advise to shifters:

- please restart supermon per run - this will keep scaler plots valid
- please do an autoprint at end of each run.
- please print the e-log from the shift and save in 'binder'



- high field run should only be done after the PS taps are changed.




   
----------------------------------------------------------------
Flemming Videbaek
Physics Department
Brookhaven National Laboratory

e-mail: videbaek@bnl.gov
phone: 631-344-4106



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 16 16:44:10 2005

This archive was generated by hypermail 2.1.8 : Sun Jan 16 2005 - 16:44:33 EST